OpenStudy (anonymous):

You invest $6000 in two account paying 6% and 9% annual interest. At the end of the year, the account earn the same interest. How much was invest at each rate?

OpenStudy (anonymous):

P(1+.06) = P(1+.09) solve

OpenStudy (anonymous):

x + y = 6000 .06x = .09y

OpenStudy (anonymous):

x + y = 6000 .06x - .09y = 0 (multiply by 100)

OpenStudy (anonymous):

x + y = 6000 6x - 9y = 0

OpenStudy (anonymous):

9(x + y = 6000) 6x - 9y = 0 9x + 9y = 54000 6x - 9y = 0

OpenStudy (anonymous):

15x = 54000 x = 3600

OpenStudy (anonymous):

y = 2400

OpenStudy (anonymous):

So $3600 at 6% and $2400 at 9%

OpenStudy (anonymous):

(Thanks, blexting and indy25, I also have another problem) The manager of a Starbuck store plans to mix A-grade coffee that cost $9.50per pound with B-grade coffee that costs $7.00per pound to create a 20-pound blend that will sell for $8.50 per pound. How many pounds of each type of coffee are required?

OpenStudy (anonymous):

lol, I would put in 20 lbs of $7/pound